# Polaritonic control of blackbody infrared radiative dissociation

### v1

@@ III.1 Microcavity-assisted blackbody infrared radiative dissociation
Equipped with the knowledge of the most relevant transitions for BIRD of HF and LiH, we can readily explain the microscopic origin of the intricate pattern in Fig. 2 by considering how the microcavity DOS changes around the previously mentioned overtones. Fig. 3 presents the ratio of the microcavity DOS to the free space for selected values of $L_{C}$ where BIRD suppression or enhancement was verified in Fig. 2 (red circles). Overtone frequencies are highlighted with dotted vertical lines, color-coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity. In Fig. 3a ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig. 2. In contrast, when the microcavity length is slightly increased to $L_{C}=3.21$ $\mu$m (Fig. 3(b) the photon DOS is enhanced at $\omega$ corresponding to the $16\rightarrow 23$ and $19\rightarrow 23$ transitions, and we observe a corresponding leap in the BIRD rate in Fig. 2. Similarly, in Fig. 3(c) ($L_{C}=9.4~{}\mu\text{m})$, all of the relevant LiH overtones but $21\rightarrow 29$ are suppressed, leading to a reduced BIRD rate in this microcavity relative to free space, whereas Fig. 3(d) shows that at $L_{C}=9.6~{}\mu\text{m}$ most of the relevant LiH overtones are enhanced, thus explaining the sudden increase in BIRD rate at $L_{C}=9.6~{}\mu\text{m}$ relative to $L_{C}=9.4~{}\mu\text{m}$.
@@
As we conclude this subsection, we note that suppression effects are nearly negligible in Fig. 2 because the dissociative process has many pathways. Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always nearly negligible as the molecular system explores alternative dissociation pathways. Therefore, our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential. Strictly speaking, we have shown that these conditions rigorously apply to diatomic BIRD, but they are likely to be even more relevant in polyatomic processes.
III.2 Polariton-assisted blackbody infrared radiative dissociation
@@
To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes. Figure 1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photonfrequency-momentum dispersion relation [Fig. 1 (e)] and the electromagnetic DOS. We explore the implications for diatomic BIRD below.
Figure 4 presents the ratio of polariton-assisted HF and LiH BIRD rates to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$. These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 at microcavity length $L_{C}$ fixed at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH, respectively. As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime. We observe prominent peaks for both HF and LiH where the relative dissociation rate is notably enhanced. Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$, where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH. Conversely, as seen in the inset of Figs. 4(a) and (c), when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed. This happens because when $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the relevant overtone transitions fall into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig. 1e). Nonetheless, the suppression is small since, as discussed above, BIRD may proceed via other pathways.
The effect of a variation of collective coupling strengths ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig. 4(b) and (d). When $\omega_{M}-\omega_{i\rightarrow j}$ is around 1-10 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates. However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively. This plateau appears to be insensitive to the magnitude of the Rabi splitting. It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig. 5.
@@
Figure 5: Photon-weighted polariton DOS $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at $\omega_{M}=1401$ cm-1. The curves obtained with distinct $\Omega_{R}$ values illustrate its effect on the photon-weighted polariton DOS relevant for radiative processes mediated by polaritonic systems.
@@
In Fig. 5, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=9.6$ $\mu$m and $\omega_{M}=v_{22\rightarrow 29}(\text{LiH})=1401$ cm-1. The case where $\Omega_{R}=0$ represents the weak coupling limit and is included here to directly compare $D_{P}(\omega)$ with $D_{C}(\omega)$. It can be seen from Fig. 5, the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap. Most remaining features, e.g., the oscillations in $D_{P}(\omega)$ with $\omega$, are also observed in weak coupling and can be understood from Eq. 8. The different curves in Fig. 5 show how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ grows with increasing $\Omega_{R}$. Therefore, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting. This becomes more likely for larger values of $\Omega_{R}$.
@@
While the singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate, our results demonstrate that this singularity does not lead to unboundedly fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting. Therefore, our results obtained as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here. In experiments, imperfections in the electromagnetic device and the strongly coupled material smooth out the photon-weighted DOS near the stopgap. Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder Satanin et al. (2005). The introduction of a UV cutoff and loss pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig. 4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc.). At the same time, the reported theoretical upper bounds on diatomic BIRD rates (Fig. 4) remain valid. Finally, we emphasize that the same qualitative behavior is observed for polariton-mediated BIRD in LiH and HF. Since these chemical species have significantly different energetics, we expect the main trends reported here also apply to other diatomic and polyatomic molecules.
@@ IV Conclusions
This work presented a detailed analysis of BIRD in infrared microcavities. Our analysis revealed mechanisms for controlling dissociation phenomena in microcavities, particularly by manipulating the electromagnetic density of states (DOS) in weak and strong light-matter coupling regimes. Overtones play a critical role in the dissociation process under both weak and strong coupling cases, as the corresponding transitions act as rate-limiting steps for the dissociative process. Therefore, the radiative dissociation rate changes induced by a microcavity are essentially tied to how the photonic density of states is modified around specific overtone transitions.
We reported microcavity-assisted dissociation rate enhancements of $\mathcal{O}(1)$ in the weak coupling regime. In microcavities where $L_{C}\leq 5~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates. As $L_{C}$ increases beyond a few microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones. Further increases in $L_{C}$ lead to BIRD rates approaching the free-space limit.
In the strong coupling regime, where polaritons mediate BIRD, we reported $\mathcal{O}(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process. This enhancement is due to the emergence of a singularity in the photon-weighted polariton DOS, which massively increases transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below. The singularity is attributed to the lack of a high-energy cutoff for the electromagnetic field and the lossless character of the strongly coupled system. Nevertheless, while the singular behavior of the photon-weighted polariton DOS is absent from experiments, we have argued its effects are expected to persist to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder. Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced changes in diatomic thermal radiative dissociation.
Our work focused on the dissociation of diatomic molecules for simplicity. The main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly and are expected to be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.
Acknowledgments. R.F.R. acknowledges support from the ACS Petroleum Research Fund under project PRF-66399-DNI6, NSF CAREER award Grant No. CHE-2340746, and startup funds provided by Emory University.

### v2

@@ III   Results and Discussion
Equipped with the knowledge of the most relevant transitions for the investigated BIRD of HF and LiH via the most energetic bound state, we can readily explain the microscopic origin of the intricate pattern in Fig. 2 by considering how the microcavity DOS changes around the previously mentioned overtones. Fig. 3 presents the ratio of the microcavity DOS to the free space for selected values of $L_{C}$ where BIRD suppression or enhancement was verified in Fig. 2 (red circles). Overtone frequencies are highlighted with dotted vertical lines, color-coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity. In Fig. 3(a) ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig. 2. In contrast, when the microcavity length slightly increases to $L_{C}=3.21$ $\mu$m [Fig. 3(b)] the photon DOS is enhanced at the $16\rightarrow 23$ and $19\rightarrow 23$ transitions frequencies, and we observe a corresponding leap in the BIRD rate in Fig. 2. Similarly, in Fig. 3(c) ($L_{C}=9.4~{}\mu\text{m})$, all of the relevant LiH overtones but $21\rightarrow 29$ are suppressed, leading to a reduced BIRD rate in this microcavity relative to free space, whereas Fig. 3(d) shows that at $L_{C}=9.6~{}\mu\text{m}$ most of the relevant LiH overtones are enhanced, thus explaining the sudden increase in BIRD rate at $L_{C}=9.6~{}\mu\text{m}$ relative to $L_{C}=9.4~{}\mu\text{m}$. Introducing a finite linewidth to the considered radiative transitions would smooth out the oscillations observed in Fig. 2 and dampen their amplitude. Consequently, the weakly coupled microcavity effect on BIRD would be significant only at lengths less than a few microns.
@@
Note that suppression effects are generally minor in Fig. 2 because the dissociative process has many pathways. Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always negligible as the molecular system explores alternative dissociation pathways. Our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
Polariton-assisted blackbody infrared radiative dissociation
@@
To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes. Fig. 1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photon frequency-momentum dispersion relation [Fig. 1 (e)] and the electromagnetic DOS. We explore the implications for diatomic BIRD via the most energetic bound state below.
@@
Fig. 4 presents the ratio of polariton-assisted BIRD to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$. These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 by fixing $L_{C}$ at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH respectively. As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime. We observe prominent peaks for HF and LiH where the relative dissociation rate is notably enhanced. Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$. Here, we focus on the case where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH. In contrast, the inset of Figs. 4(a) and (c) show that when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed. This suppression occurs because as $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the overtone transition falls into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig. 1e). However, the BIRD suppression is relatively small since, as discussed earlier, the reaction may still proceed through alternative pathways.
The effect of the collective coupling strength ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig. 4(b) and (d). When $\omega_{M}-\omega_{i\rightarrow j}$ is around 0.01-1 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates. However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively. This plateau appears to be insensitive to the magnitude of the Rabi splitting. It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig. 5.
Figure 5: Photon-weighted polariton density of states $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with length $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at frequency $\omega_{M}=1401$ cm-1. The curves show the effect of different Rabi splitting values ($\Omega_{R}$) on the photon-weighted polariton DOS, which is relevant for radiative processes mediated by polaritonic systems. As $\Omega_{R}$ increases, the frequency range below $\omega_{M}$, where the photon-weighted polariton DOS significantly exceeds that of free space, widens. This trend explains the impact of the collective light-matter interaction strength on the observed polariton-assisted bond infrared dissociation (BIRD) enhancement shown in Fig. 4.
@@
In Fig. 5, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=9.6$ $\mu$m and $\omega_{M}=v_{22\rightarrow 29}(\text{LiH})=1401$ cm-1. The case where $\Omega_{R}=0$ represents the weak coupling limit and is included here to directly compare $D_{P}(\omega)$ with $D_{C}(\omega)$. It can be seen from Fig. 5 that the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap. Most remaining features, e.g., the oscillations in $D_{P}(\omega)$ with $\omega$, are also observed in weak coupling and can be understood from Eq. 8. The various curves in Fig. 5 illustrate how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ expand as $\Omega_{R}$ increases. As a result, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting.
@@
The singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate. However, our results demonstrate that this singularity does not lead to arbitrarily fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting. Therefore, the BIRD rates reported as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here. In experiments, imperfections in the electromagnetic device and the strongly coupled material smooth out the photon-weighted DOS near the stopgap. Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder [63]. The introduction of a UV cutoff, loss, and dephasing pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig. 4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc).
Including a finite linewidth for the diatomic vibrational transitions relaxes the strict resonance requirement discussed above ($\omega_{M}\rightarrow\omega_{i\rightarrow j}$), and the overall enhancement will depend on the spectral overlap between $\omega_{i\rightarrow j}$ and $D_{P}(\omega)$. Since the molecular linewidth and cavity imperfections are expected to be limiting factors, it is clear that our reported results represent theoretical upper bonds for polariton-assisted BIRD rates for the reaction mechanism investigated here.
Before ending, we emphasize our observation of the same qualitative behavior for polariton-mediated BIRD in LiH and HF. Since these chemical species have significantly different energetics, we anticipate the main trends reported in this work also apply to other diatomic and polyatomic molecules, particularly in cases where dissociation occurs through a specific doorway state.
@@ IV   Conclusions
This work presented a detailed analysis of BIRD in infrared microcavities. Our analysis revealed mechanisms for controlling dissociation phenomena in microcavities, particularly by manipulating the electromagnetic density of states (DOS) in weak and strong light-matter coupling regimes. Overtones play a critical role in the dissociation process under both weak and strong coupling cases, as the corresponding transitions act as rate-limiting steps for the dissociative process occurring via the highest energy-bound state. The radiative dissociation rate changes induced by a microcavity are essentially tied to how the photonic density of states is modified around specific overtone transitions driving the system into the dissociating (doorway) state.
We reported microcavity-assisted dissociation rate enhancements of $O(1)$ in the weak coupling regime. In microcavities where $L_{C}\leq 5~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates. As $L_{C}$ increases beyond a few microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones. Further increases in $L_{C}$ causes the BIRD rates to approach the free-space limit.
In the strong coupling regime, where polaritons mediate BIRD, we reported $O(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process. This enhancement is due to the emergence of a singularity in the photon-weighted polariton DOS, which massively increases transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below. The singularity is attributed to the lack of a high-energy cutoff for the electromagnetic field and the lossless character of the strongly coupled system. Nevertheless, while the singular behavior of the photon-weighted polariton DOS is absent from experiments, we have argued its effects are expected to persist to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder. Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced diatomic thermal radiative dissociation changes via the most energetic bound state.
Our work focused on how optical microcavities can modify BIRD reactions of diatomic molecules via a particular mechanism. Despite focusing on these specific systems, the main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly. We expect the analysis introduced here will be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.
Acknowledgments. This work was supported by the donors of ACS Petroleum Research Fund under Doctoral New Investigator Grant 66399-DNI6. R.F.R. served as Principal Investigator on ACS PRF-66399-DNI6 that provided support for E. S. R.F.R also acknowledges support from NSF CAREER award Grant No. CHE-2340746, and startup funds provided by Emory University.